OpenStudy (anonymous):

Can someone please help me? I've been having a lot of trouble on this, and now this assignment is a week overdue. I just need it explained to me, on how to do it, then I can do it on my own.

OpenStudy (anonymous):

Number 6 is the problem I'm on, if I remember correctly.

OpenStudy (anonymous):

Actually it's number 9. Below it.

OpenStudy (kropot72):

 This is the graph of y = |x| To translate the graph to y = |x| - 3 each value of y needs to have 3 subtracted from it. Can you see in what direction the graph will move if this subtraction is done?

OpenStudy (phi):

You really don't have to think too hard to do this. notice that y= |x| makes a "V" with the vertex at x=0. for y= |x|-3 replace x with 0 and you find y= |0|-3 = -3 so one point is going to be 0,-3 now use x=1 and you find y= -2, so another point is (1,-2) plot those two points and draw a line through them now use x=-1, and you find y= -2 (again). plot the point (-1,-2) draw a line through the points (-1,-2) and (0,-3) now "clean it up" so that you get a "V" shaped graph

OpenStudy (anonymous):

Thanks both of you guys, math really isn't my strong point.

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

y= |x| is the parent function any changes such as y= |x| -3 chang where it is graphed.. Just to add to what they have said :)

OpenStudy (anonymous):

Thanks everyone. :D

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

@kropot72 @phi @Shadowthieve how would I go about doing numbers 17 - 20, I'm a bit confused on those.

OpenStudy (anonymous):

@kropot72 @phi @Shadowthieve And the ones after that as well, these ones confuse me.

OpenStudy (phi):

For question 17, 0.25 units up First, notice that each "box" has unit length (each side is length 1) they want you to write an equation that moves the graph y = |x| up 1/4 of a box y=|x| at 0 has a y value of y= |0| (replace x with a zero), you get y=0 if the graph were up 0.25 (1/4 of a box), y would be 0.25 so maybe you can see that the equation you want is y = |x| +0.25 all the other problems 18-20 are the same, except down means add a minus number (and of course "up" means add a positive number) to do problems 18-20, add the number they give you (instead of 0.25). and make the number minus if you want "down"

OpenStudy (phi):

For problem 21, find the x value that makes y=0 (this will be where the vertex (V part) is) y = | x -3 | what x makes x-3 = 0 ? add +3 to both sides and you get x-3+3= +3 simplify to get x+0= 3 or x=3 x=3 is where the V part of the graph is. The graph has move "to the right" 3 boxes.

OpenStudy (anonymous):

okay that makes a lot of sense, thank you.